# Optimal Computation in Anonymous Dynamic Networks

### v8

@@ 5.3 Subroutine ApproxCount
[…]
@@
[…]
@@
Observe that all the terms in this chain of inequalities are between the two consecutive integers $w(v)-1$ and $w(v)$. It follows that
@@
Let $s$ be a counted node in $\mathcal{V}$, and let $F$ be a cut for $v$ whose nodes are all counted. Then, the set of nodes spanned by the black paths from $s$ to the nodes of $F$ is called isle; $s$ is the root of the isle, while each node in $F$ is a leaf of the isle. The nodes in an isle other than the root and the leaves are called internal. An isle is said to be trivial if it has no internal nodes.
@@
[…]
@@ 18 Let $v^{\prime}$ be the heavy node in $P_{v}$ of maximum depth
21 If $v^{\prime}$ is the root or a leaf of a non-trivial complete isle $I$
22 For each internal node $w$ of $I$
@@ 24 Mark $w$ as counted, not guessed, and not locked
[…]
@@ 29 Assign $\ell^{\prime}:=\sum_{v\text{ leader node in }C}a^{\prime}(v)$
[…]
@@ 32 Else return $(n^{\prime},t)$
Assume that $\delta\geq 1$. Then, as long as level $L_{t}$ of $\mathcal{V}$ is not missing any nodes (where $t$ is defined and updated as in ApproxCount), whenever Line 13 is reached, there are at most $n-2$ levels in the range from $L_{s+1}$ to $L_{t}$ where all guessable nodes are already counted.
@@
We have proved that every level between $L_{s+1}$ and $L_{t}$ where all guessable nodes are counted contains a counted node $v$ having a parent or a sibling that is not counted: we call such a node $v$ a bad node. To conclude the proof, it suffices to show that there are at most $n-2$ bad nodes between $L_{s+1}$ and $L_{t}$. Observe that no nodes in $P_{\vartheta}$ can be bad.
We will prove that, if a subtree $\mathcal{W}$ of $\mathcal{V}$ contains the root $r$, the leader node $\vartheta$, no counting cuts, and no non-trivial isles, then $\mathcal{W}$ contains at most $f-1$ bad nodes, where $f$ is the number of leaves of $\mathcal{W}$ not in the subtree hanging from $\vartheta$. We stress that, in the context of $\mathcal{W}$, a bad node is defined as a counted node in $\mathcal{W}$ (other than $\vartheta$ and its descendants) that has a non-counted parent or a non-counted sibling in $\mathcal{W}$.
@@
For the induction step, let $v$ be a bad node of maximum depth in $\mathcal{W}$. Let $(v_{1},v_{2},\dots,v_{k})$ be the black path from $v_{1}=v$ to the root $v_{k}=r$, and let $1<i\leq k$ be the smallest index such that $v_{i}$ has more than one child in $\mathcal{W}$ ($i$ must exist, because this path eventually joins the black path from $\vartheta$ to $r$). Let $\mathcal{W}^{\prime}$ be the tree obtained by deleting the black edge $\{v_{i-1},v_{i}\}$ from $\mathcal{W}$, as well as the subtree hanging from it.
@@
Notice that the induction hypothesis applies to $\mathcal{W}^{\prime}$: since $v_{1}$ is counted, and each of the nodes $v_{2}$, …, $v_{i-1}$ has a unique child in $\mathcal{W}$, the removal of $\{v_{i-1},v_{i}\}$ does not create counting cuts or non-trivial isles. Also, $v_{2}$ is not counted (unless $v_{2}=v_{i}$), because $v_{1}$ is bad. Furthermore, none of the nodes $v_{3}$, …, $v_{i-1}$ is counted, or else $v_{2}$ would be an internal node of a (non-trivial) isle in $\mathcal{W}$. Therefore, none of the nodes $v_{2}$, …, $v_{i-1}$ is counted. In particular, none of these nodes is bad in $\mathcal{W}$.
@@
Moreover, a node of $\mathcal{W}^{\prime}$ is bad if and only if it is bad in $\mathcal{W}$. This is trivial for all nodes, except for the siblings of $v_{i-1}$, which require a careful proof. Let $u\neq v_{i-1}$ be a sibling of $v_{i-1}$ in $\mathcal{W}$. If $u$ is not counted, then it is not a bad node in $\mathcal{W}$ nor in $\mathcal{W}^{\prime}$. Hence, let us assume that $u$ is counted. If $v_{i}$ is not counted, then $u$ is a bad node in $\mathcal{W}$ and in $\mathcal{W}^{\prime}$. Thus, let us assume that $v_{i}$ is counted. Note that $i=2$ must hold, otherwise $v$ and $v_{i}$ would form a non-trivial isle in $\mathcal{W}$. In particular, $v$ and $u$ are siblings. Since $v$ is a bad node and its parent $v_{2}$ is counted, it must have a non-counted sibling $u^{\prime}\neq u$ in $\mathcal{W}$. However, $u^{\prime}$ is also a node of $\mathcal{W}^{\prime}$, and therefore $u$ is not a bad node in $\mathcal{W}$ nor in $\mathcal{W}^{\prime}$.
@@
It follows that $\mathcal{W}^{\prime}$ has exactly one less bad node than $\mathcal{W}$ and at most $f-1$ leaves (because there is at least one leaf in the subtree of $\mathcal{W}$ hanging from $v$). Thus, the induction hypothesis implies that $\mathcal{W}^{\prime}$ contains at most $f-2$ bad nodes, and therefore $\mathcal{W}$ contains at most $f-1$ bad nodes.
@@
Observe that the subtree $\mathcal{V}^{\prime}$ of $\mathcal{V}$ formed by all levels up to $L_{t}$ satisfies all of the above conditions, as it contains $\vartheta\in L_{s}$, the root $r$, and has no counting cuts, because a counting cut for $\mathcal{V}^{\prime}$ would be a counting cut for $\mathcal{V}$, as well (recall that $\mathcal{V}$ has no counting cuts). Also, Lemma 5.6 (iii) ensures that $\mathcal{V}^{\prime}$ contains no non-trivial complete isles. However, since no nodes are missing from the levels of $\mathcal{V}^{\prime}$, all isles in $\mathcal{V}^{\prime}$ are complete, and thus must be trivial. We conclude that, if $\mathcal{V}^{\prime}$ has $f$ leaves not in the subtree hanging from $\vartheta$, it contains at most $f-1$ bad nodes. Since such leaves induce a partition of the at most $n-1$ agents not represented by $\vartheta$, we have $f\leq n-1$, implying that the number of bad nodes up to $L_{t}$ is at most $n-2$. ∎
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
Worst-case example. For $\ell=\tau=1$, our Counting algorithm in Theorem 5.2 yields a running time of $23n-$ 2 rounds. The example in Figure 8, which can easily be generalized to networks of any size $n$, shows that our Counting algorithm can in fact terminate in $33n-$ 3 rounds. Indeed, the last node in a counting cut is at level $L_{2n-3}$, and then it takes an extra $n$ rounds for the terminating condition $t\geq t^{\ast}+n^{\ast}$ of Listing 2 to be satisfied.
Figure 8: An example of a dynamic network with $\ell=\tau=1$ and $n=6$ where the algorithm of Theorem 5.2 terminates in $33n-$ 3 rounds, almost matching the upper bound of $23n-$ 2 rounds. The white nodes in the history tree are not in the vista of the leader at the last round. The numbers inside nodes represent guesses; the blue ones correspond to nodes that are marked as counted. The six highlighted nodes constitute a counting cut.
@@ 6.1 Unsolvable Problems
Networks with leaders. We will prove that the multiset-based functions introduced in Section 2 are the only functions that can be computed deterministically in anonymous networks with leaders, even when restricted to connected static simply networks. This result, together with Proposition 2.1, shows that the algorithms in Sections 4.4 and 5 can be generalized to all functions that can be computed in networks with leaders, i.e., the multiset-based functions.
@@
For any $\ell\geq 1$, no algorithm computes a non-multiset-based function (with or without termination) on all static simple networks in $\bigcup_{n\geq\ell}\mathcal{N}^{1}_{n,\ell}$.
@@
For $n\geq\ell$, let us consider the (static) network whose topology at round $t$ is the complete graph $G_{t}=K_{n}$, i.e., each agent receives messages from all other agents at every round. We can prove by induction that all nodes of the history tree other than the root have exactly one child. This is because any two agents with the same input always receive equal multisets of messages, and are therefore always indistinguishable. Thus, the history tree is completely determined by the multiset $\mu_{\lambda}$ of all agents’ inputs; moreover, an agent’s vista at any given round only depends on the agent’s own input and on $\mu_{\lambda}$ (the fact that $\ell$ is a constant does not affect this). By the fundamental theorem of history trees Theorem 3.1, this is enough to conclude that if an agent’s output stabilizes, that output must be a function of the agent’s own input and of $\mu_{\lambda}$, which is the defining condition of a multiset-based function. ∎
The following result justifies the assumption made in Sections 4.4 and 5 that agents have a-priori knowledge of the number of leaders $\ell$ in the system. It states that no algorithm can compute the Counting function $F_{C}$ (which, as we recall, is a multiset-based function) with no exact knowledge of $\ell$, even when restricted to connected static simple networks with a known and arbitrarily small ratio $\ell/n$.
@@
For any integer $k\geq 1$, no algorithm computes the Counting function $F_{C}$ (with or without termination) on all static simple networks in $\bigcup_{\ell\geq 1}\mathcal{N}^{1}_{k\ell,\ell}$.
@@
Let us fix $k\geq 1$; we will construct an infinite class of networks with $k\ell$ agents and $\ell$ leaders as follows. For every $i\geq 3$, let $\mathcal{G}_{i}$ be the static network consisting of a cycle of $n_{i}=k\cdot i$ agents of which $\ell_{i}=i$ are leaders, such that the leaders are evenly spaced among the non-leaders. Assume that all agents get the same input (apart from their leader flags). Then, at any round, all the leaders in all of these networks have isomorphic vistas, which are independent of $i$. It follows that, since nothing is known about $\ell_{i}$ (other than the ratio $\ell_{i}/n_{i}$, which is fixed), all the leaders in all the networks $\mathcal{G}_{i}$ always give equal outputs. Since the number of agents $n_{i}$ depends on $i$, at most one of these networks can stabilize on the correct output $n_{i}$. ∎
@@
Let $m_{1}$, $m_{2}$, …, $m_{k}$ be integers greater than $2$ with $\gcd(m_{1},m_{2},\dots,m_{k})=1$, and let $B$ be the complete $k$-partite graph with partite sets $V_{1}$, $V_{2}$, …, $V_{k}$ of sizes $m_{1}$, $m_{2}$, …, $m_{k}$, respectively. For any positive integer $\alpha$, construct the static network $G_{\alpha}$ consisting of $\alpha$ disjoint copies of $B$, augmented with $k$ cycles $C_{1}$, $C_{2}$, …, $C_{k}$ such that, for each $1\leq i\leq k$, the cycle $C_{i}$ spans all the $\alpha m_{i}$ agents in the $\alpha$ copies of $V_{i}$. Clearly, $G_{\alpha}$ is a connected static simple network.
Let the function $\lambda_{\alpha}$ assign input $z_{i}$ to all agents in the $\alpha$ copies of $V_{i}$ in $G_{\alpha}$, and let $\lambda=\lambda_{1}$. As a result, $\mu_{\lambda_{\alpha}}=\{(z_{1},\alpha m_{1}),(z_{2},\alpha m_{2}),\dots,(z_{k},\alpha m_{k})\}=\alpha\cdot\mu_{\lambda}$ for all $\alpha\geq 1$. Moreover, all the networks $G_{\alpha}$ have isomorphic history trees. This is because, at every round, each agent in any of the copies of $V_{i}$ receives exactly two messages from other agents in copies of $V_{i}$ and exactly $m_{j}$ messages from agents in copies of $V_{j}$, for all $j\neq i$. Thus, it can be proved by induction that all agents in the copies of $V_{i}$ have isomorphic vistas, regardless of $\alpha$.
Due to the fundamental theorem of history trees Theorem 3.1, all the agents with input $z_{i}$ must give the same output $\psi(z_{i},\mu_{\lambda})=\psi(z_{i},\mu_{\lambda_{\alpha}})=\psi(z_{i},\alpha\cdot\mu_{\lambda})$, regardless of $\alpha$. Hence, by definition, only frequency-based functions can be computed in these networks. ∎
@@
Assume for a contradiction that there is such an algorithm $\mathcal{A}$. Let $\mathcal{G}$ be a static network consisting of three agents forming a cycle, and assign input 00 to all of them. If the agents execute $\mathcal{A}$, they eventually output the mean value 00 and terminate, say in $t$ rounds.
Now construct a static network $\mathcal{G}^{\prime}$ consisting of a cycle of $22t+$ 2 agents $p_{1}$, $p_{2}$, …, $p_{2t+2}$; assign input $1$ to $p_{1}$ and input 00 to all other agents. It is easy to see that, from round 00 to round $t$, the vista of the agent $p_{t+1}$ is isomorphic to the vista of any agent in $\mathcal{G}$. Therefore, if $p_{t+1}$ executes $\mathcal{A}$, it terminates in $t$ rounds with the incorrect output 00. Thus, $\mathcal{A}$ is incorrect. ∎

### v9

@@ 5.3 Subroutine ApproxCount
Overview. In ApproxCount, we are given a vista $\mathcal{V}$ with a strand of leader nodes hanging from the first leader node $\vartheta$ in level $L_{s}$, where the anonymity $a(\vartheta)$ is an unknown number not greater than $\ell$. The algorithm begins by assuming that $a(\vartheta)$ is the given parameter $x$, and then it makes deductions on the anonymities of other nodes until it is able to either make an estimate $n^{\prime}>0$ on the total number of agents or report failure in the form of an error message $n^{\prime}\in\{\texttt{"MissingNodes"},\texttt{"StrandTooShort"},\texttt{"WrongGuess"}\}$. In particular, since the algorithm requires the existence of a long-enough strand hanging from $\vartheta$, it reports failure if some descendants of $\vartheta$ (in the relevant levels of $\mathcal{V}$) have more than one child.
@@
Discrepancy $\delta$. ApproxCount is invoked with arguments $\mathcal{V}$, $s$, $x$, $\ell$, where $1\leq x\leq\ell$. If there are no nodes representing leaders in level $L_{s}$ of $\mathcal{V}$, the procedure immediately returns the error message $n^{\prime}=\texttt{"MissingNodes"}$. Otherwise, we denote by $\vartheta$ the first such node. We define the discrepancy $\delta$ as the ratio $x/a(\vartheta)$. Since $a(\vartheta)$ is not known in advance by the agent executing ApproxCount, then neither is $\delta$. Nonetheless, since both $x$ and $a(\vartheta)$ range between 11 and $\ell$, it follows that $1/\ell\leq\delta\leq\ell$.
@@
Observe that all the terms in this chain of inequalities are between the two consecutive integers w(v)1w(v)-1 and $w(v)$. It follows that
@@
Let $s$ be a counted node in $\mathcal{V}$, and let $F$ be a cut for $s$ whose nodes are all counted. Then, the set of nodes spanned by the black paths from $s$ to the nodes of $F$ is called isle; $s$ is the root of the isle, while each node in $F$ is a leaf of the isle. The nodes in an isle other than the root and the leaves are called internal. An isle is said to be trivial if it has no internal nodes.
@@
6If $L_{s}$ does not contain any leader nodes, return $(\texttt{"MissingNodes"},t)$
@@ 19 If $g(v^{\prime})$ is not an integer, return $(\texttt{"WrongGuess"},t)$
21 If $v^{\prime}$ is the root or a leaf of a non-trivial complete isle II
22 For each internal node $w$ of II
@@ 24 Mark $w$ as counted, not guessed, and not locked
25If no counting cut has been found, return $(\texttt{"StrandTooShort"},t)$
@@ 29 Assign $\ell^{\prime}:=\sum_{v\text{ leader node in }C}a^{\prime}(v)$
30 If $\ell^{\prime}<\ell$, return $(\texttt{"MissingNodes"},t)$
31 Else if $\ell^{\prime}>\ell$, return $(\texttt{"WrongGuess"},t)$
@@ 32 Else return $(n^{\prime},t)$
Assume that $\delta\geq 1$. Then, as long as level $L_{t}$ of $\mathcal{V}$ is not missing any nodes (where $t$ is defined and updated as in ApproxCount), whenever Line 13 is reached, there are at most n2n-2 levels in the range from $L_{s+1}$ to $L_{t}$ where all guessable nodes are already counted.
@@
We have proved that every level between $L_{s+1}$ and $L_{t}$ where all guessable nodes are counted contains a counted node $v$ having a parent or a sibling that is not counted: we call such a node $v$ a bad node. To conclude the proof, it suffices to show that there are at most n2n-2 bad nodes between $L_{s+1}$ and $L_{t}$. Observe that no nodes in $P_{\vartheta}$ can be bad.
We will prove that, if a subtree $\mathcal{W}$ of $\mathcal{V}$ contains the root $r$, the leader node $\vartheta$, no counting cuts, and no non-trivial isles, then $\mathcal{W}$ contains at most f1f-1 bad nodes, where $f$ is the number of leaves of $\mathcal{W}$ not in the subtree hanging from $\vartheta$. We stress that, in the context of $\mathcal{W}$, a bad node is defined as a counted node in $\mathcal{W}$ (other than $\vartheta$ and its descendants) that has a non-counted parent or a non-counted sibling in $\mathcal{W}$.
@@
For the induction step, let $v$ be a bad node of maximum depth in $\mathcal{W}$. Let $(v_{1},v_{2},\dots,v_{k})$ be the black path from $v_{1}=v$ to the root $v_{k}=r$, and let $1<i\leq k$ be the smallest index such that $v_{i}$ has more than one child in $\mathcal{W}$ (ii must exist, because this path eventually joins the black path from $\vartheta$ to $r$). Let $\mathcal{W}^{\prime}$ be the tree obtained by deleting the black edge $\{v_{i-1},v_{i}\}$ from $\mathcal{W}$, as well as the subtree hanging from it.
@@
Note that the induction hypothesis applies to $\mathcal{W}^{\prime}$: since $v_{1}$ is counted, and each of the nodes $v_{2}$, …, $v_{i-1}$ has a unique child in $\mathcal{W}$, the removal of $\{v_{i-1},v_{i}\}$ does not create counting cuts or non-trivial isles. Also, $v_{2}$ is not counted (unless $v_{2}=v_{i}$), because $v_{1}$ is bad. Furthermore, none of the nodes $v_{3}$, …, $v_{i-1}$ is counted, or else $v_{2}$ would be an internal node of a (non-trivial) isle in $\mathcal{W}$. Therefore, none of the nodes $v_{2}$, …, $v_{i-1}$ is counted. In particular, none of these nodes is bad in $\mathcal{W}$.
@@
Moreover, a node of $\mathcal{W}^{\prime}$ is bad if and only if it is bad in $\mathcal{W}$. This is trivial for all nodes, except for the siblings of $v_{i-1}$, which require a careful proof. Let $u\neq v_{i-1}$ be a sibling of $v_{i-1}$ in $\mathcal{W}$. If $u$ is not counted, then it is not a bad node in $\mathcal{W}$ nor in $\mathcal{W}^{\prime}$. Hence, let us assume that $u$ is counted. If $v_{i}$ is not counted, then $u$ is a bad node in $\mathcal{W}$ and in $\mathcal{W}^{\prime}$. Thus, let us assume that $v_{i}$ is counted.
Assume for a contradiction that all children of $v_{i}$ other than $v_{i-1}$ are counted. Then, if $i>2$, the nodes $v$, $v_{i}$, and the children of $v_{i}$ (other than $v_{i-1}$) are all counted and form a non-trivial isle in $\mathcal{W}$, which is impossible. On the other hand, if $i=2$, then $v_{i}$ and all its children (including $v=v_{1}$) are counted, which contradicts the fact that $v$ is a bad node. Thus, we have proved that $v_{i}$ must have a non-counted child in $\mathcal{W}^{\prime}$ (hence in $\mathcal{W}$), and therefore $u$ is not a bad node in $\mathcal{W}$ nor in $\mathcal{W}^{\prime}$.
@@
It follows that $\mathcal{W}^{\prime}$ has exactly one less bad node than $\mathcal{W}$ and at most f1f-1 leaves (because there is at least one leaf in the subtree of $\mathcal{W}$ hanging from $v$). Thus, the induction hypothesis implies that $\mathcal{W}^{\prime}$ contains at most f2f-2 bad nodes, and therefore $\mathcal{W}$ contains at most f1f-1 bad nodes.
@@
Observe that the subtree $\mathcal{V}^{\prime}$ of $\mathcal{V}$ formed by all levels up to $L_{t}$ satisfies all of the above conditions, as it contains $\vartheta\in L_{s}$, the root $r$, and has no counting cuts, because a counting cut for $\mathcal{V}^{\prime}$ would be a counting cut for $\mathcal{V}$, as well (recall that $\mathcal{V}$ has no counting cuts). Also, Lemma 5.6 (iii) ensures that $\mathcal{V}^{\prime}$ contains no non-trivial complete isles. However, since no nodes are missing from the levels of $\mathcal{V}^{\prime}$, all isles in $\mathcal{V}^{\prime}$ are complete, and thus must be trivial. We conclude that, if $\mathcal{V}^{\prime}$ has $f$ leaves not in the subtree hanging from $\vartheta$, it contains at most f1f-1 bad nodes. Since such leaves induce a partition of the at most n1n-1 agents not represented by $\vartheta$, we have $f\leq n-1$, implying that the number of bad nodes up to $L_{t}$ is at most n2n-2. ∎
@@
    If $x=a(\vartheta)=a(\vartheta^{\prime})$, then $n^{\prime}\neq\texttt{"WrongGuess"}$.
@@
    If $t^{\prime}\geq s+(\ell+2)(n-1)$, then $s\leq t\leq s+(\ell+1)(n-1)$ and $n^{\prime}\neq\texttt{"MissingNodes"}$. Moreover, if $n^{\prime}=\texttt{"StrandTooShort"}$, then $L_{t}$ contains a leader node with at least two children in $\mathcal{V}$.
@@
Let us prove statement (i). The assumption $x=a(\vartheta)$ implies $\delta=1$. Moreover, since $a(\vartheta)=a(\vartheta^{\prime})$, 1 is satisfied whenever a guess is made (this is a straightforward induction). Therefore, by Lemma 5.6 (iv), all nodes marked as counted up to $L_{t}$ indeed have the correct guesses. So, the conditional anonymity that is computed for any node is equal to its anonymity ($a^{\prime}(v)=\delta a(v)=a(v)$), and hence is an integer. This implies that ApproxCount cannot return the error message "WrongGuess" at Line 19. Also, either $\ell^{\prime}=\ell$ if all leader agents have been counted, or $\ell^{\prime}<\ell$ if some leader nodes are missing from the vista. Either way, ApproxCount cannot return the error message "WrongGuess" at Line 31. We conclude that $n^{\prime}\neq\texttt{"WrongGuess"}$.
@@
Assume that $p<s+(\ell+1)(n-1)$. This implies that no nodes are missing from level $L_{p+1}$, and therefore $\vartheta^{\prime\prime}$ must have at least two children in $L_{p+1}$. Since $t\leq p$, we have $t<s+(\ell+1)(n-1)$, as desired. Now assume that $n^{\prime}=\texttt{"StrandTooShort"}$, which implies that the algorithm was unable to find a counting cut. We claim that in this case $t=p$. So, assume for a contradiction that $t\leq p-1$. It follows that $\vartheta^{\prime}\in L_{t}$ is a guesser. Recall that $L_{t}$ and $L_{t+1}$ are not missing any nodes, because $t\leq p$. Since the network is connected at round $t+1$, there is at least one node $v\in L_{t+1}$ that is guessable by $\vartheta^{\prime}$. Also, since no guess has ever been made in level $L_{t+1}$ (due to the way $t$ is updated in Line 14), it follows that $v$ is not counted and not locked. However, the algorithm cannot return $n^{\prime}=\texttt{"StrandTooShort"}$ as long as there are nodes such as $v$ (Line 12). Thus, $t=p$, which means that $\vartheta^{\prime}=\vartheta^{\prime\prime}$, and hence $\vartheta^{\prime}$ has at least two children in $\mathcal{V}$, as desired.
Assume now that $p\geq s+(\ell+1)(n-1)$. If $\vartheta$ is the only node in $L_{s}$, it constitutes a counting cut. In this special case, Line 13 is never reached, $n^{\prime}=\texttt{"StrandTooShort"}$ is not returned at Line 25, and $t=s$. Hence, we may assume that there are nodes in $L_{s}$ other than $\vartheta$ and Line 13 is reached (so, we have $n>1$). Recall that no nodes are missing from any level up to $L_{s+(\ell+1)(n-1)}$. In particular, no nodes are missing from the levels in the non-empty interval $\mathcal{L}$ consisting of the $(\ell+1)(n-1)$ levels from $L_{s+1}$ to $L_{s+(\ell+1)(n-1)}$. Thus, by definition of $p$, as long as no guesses are made outside of $\mathcal{L}$, 1 holds, and therefore Lemmas 5.7 and 5.8 apply. Hence, as long as no guesses are made outside of $\mathcal{L}$, there are at most $\delta(n-1)$ levels of $\mathcal{L}$ containing locked nodes (Lemma 5.7) and there are at most n2n-2 levels of $\mathcal{L}$ where all guessable nodes are already counted (Lemma 5.8). So, there are at most
@@
levels in $\mathcal{L}$ were the algorithm can make no new guesses. We conclude that $\mathcal{L}$ always contains at least one node where a guess can be made, and no guesses are ever made outside of $\mathcal{L}$ until either a counting cut is found or $n^{\prime}=\texttt{"WrongGuess"}$ is returned at Line 19. In both cases, $n^{\prime}=\texttt{"StrandTooShort"}$ is not returned, and moreover $t\leq s+(\ell+1)(n-1)$, as desired.
It remains to prove that $n^{\prime}\neq\texttt{"MissingNodes"}$. Since $\vartheta$ exists by assumption, the error message "MissingNodes" cannot be returned at Line 6 and can only be returned at Line 30. In turn, this can only occur if a counting cut has been found and $\ell^{\prime}<\ell$. However, we have already proved that no level up to $L_{t}$ is missing any nodes, which implies that the counting cut contains nodes representing all agents, and in particular $\ell^{\prime}=\delta\ell$. Since $\delta\geq 1$, we have $\ell^{\prime}\geq\ell$, and the condition at Line 30 is not satisfied. ∎
Worst-case example. For $\ell=\tau=1$, our Counting algorithm in Theorem 5.2 yields a running time of 3n23n-2 rounds. The example in Figure 8, which can easily be generalized to networks of any size $n$, shows that our Counting algorithm can in fact terminate in 3n33n-3 rounds. Indeed, the last node in a counting cut is at level $L_{2n-3}$, and then it takes an extra $n$ rounds for the termination condition $t\geq t^{\ast}+n^{\ast}$ of Listing 2 to be satisfied.
Note that removing all the double edges from this network and eliminating round $1$ yields a simple dynamic network where our Counting algorithm terminates in 3n43n-4 rounds and stabilization occurs in 2n32n-3 rounds, providing a close-to-worst-case example for both Theorems 5.2 and 4.7.
Figure 8: An example of a dynamic network with $\ell=\tau=1$ and $n=6$ where the algorithm of Theorem 5.2 terminates in 3n33n-3 rounds, almost matching the upper bound of 3n23n-2 rounds. The white nodes in the history tree are not in the vista of the leader at the last round. The numbers inside nodes represent guesses; the blue ones correspond to nodes that are marked as counted. The six highlighted nodes constitute a counting cut.
@@ 6.1 Unsolvable Problems
Networks with leaders. We will prove that the multiset-based functions introduced in Section 2 are the only functions that can be computed deterministically in anonymous networks with leaders, even when restricted to connected static simple networks. This result, together with Proposition 2.1, shows that the algorithms in Sections 4.4 and 5 can be generalized to all functions that can be computed in networks with leaders, i.e., the multiset-based functions.
@@
For any $\ell\geq 1$ and $n\geq\ell$, no algorithm computes a non-multiset-based function (with or without termination) on all static simple networks in $\mathcal{N}^{1}_{n,\ell}$.
@@
Let us consider the (static) network whose topology at round $t$ is the complete graph $G_{t}=K_{n}$, i.e., each agent receives messages from all other agents at every round. We can prove by induction that all nodes of the history tree other than the root have exactly one child. This is because any two agents with the same input always receive equal multisets of messages, and are therefore always indistinguishable. Thus, the history tree is completely determined by the multiset $\mu_{\lambda}$ of all agents’ inputs; moreover, an agent’s vista at any given round only depends on the agent’s own input and on $\mu_{\lambda}$ (the fact that $\ell$ is a constant does not affect this). By the fundamental theorem of history trees Theorem 3.1, this is enough to conclude that if an agent’s output stabilizes, that output must be a function of the agent’s own input and of $\mu_{\lambda}$, which is the defining condition of a multiset-based function. ∎
The following result justifies the assumption made in Sections 4.4 and 5 that agents have a-priori knowledge of the number of leaders $\ell$ in the system. It states that no algorithm can compute the Counting function $F_{C}$ (which, as we recall, is a multiset-based function) with no exact knowledge of $\ell$, even when restricted to connected static simple networks with a known and arbitrarily large ratio $n/\ell=k$.
@@
For any integer $k\geq 1$ and $\ell\geq 1$, no algorithm computes the Counting function $F_{C}$ (with or without termination) on all static simple networks in $\mathcal{N}^{1}_{k\ell,\ell}\cup\mathcal{N}^{1}_{k(\ell+1),\ell+1}$.
@@
Let $\mathcal{G}$ (respectively, $\mathcal{G}^{\prime}$) be the static network consisting of a cycle of $k\ell$ (respectively, $k(\ell+1)$) agents of which $\ell$ are leaders, such that the leaders are evenly spaced among the non-leaders. That is, each leader has a sequence of exactly $k$ consecutive non-leaders on each side. Assume that all agents in $\mathcal{G}$ and $\mathcal{G}^{\prime}$ get the same input (apart from their leader flags). Then, at any round, all leaders in both networks have isomorphic vistas, and therefore always give equal outputs. However, in order to compute the Counting function, the leaders of $\mathcal{G}$ have to eventually output $k\ell$, while the leaders of $\mathcal{G}^{\prime}$ have to eventually output $k(\ell+1)$. Hence, at most one of these networks can stabilize on the correct output. ∎
@@
Let $n_{1}$, $n_{2}$, …, $n_{k}$ be integers greater than $2$ with $\gcd(n_{1},n_{2},\dots,n_{k})=1$, and let $B$ be the complete $k$-partite graph with partite sets $V_{1}$, $V_{2}$, …, $V_{k}$ of sizes $n_{1}$, $n_{2}$, …, $n_{k}$, respectively. For any positive integer $m$, construct the static network $G_{m}$ consisting of $m$ disjoint copies of $B$, augmented with $k$ cycles $C_{1}$, $C_{2}$, …, $C_{k}$ such that, for each $1\leq i\leq k$, the cycle $C_{i}$ spans all the $mn_{i}$ agents in the $m$ copies of $V_{i}$. Clearly, $G_{m}$ is a connected static simple network.
Let the function $\lambda_{m}$ assign input $z_{i}$ to all agents in the $m$ copies of $V_{i}$ in $G_{m}$, and let $\lambda=\lambda_{1}$. As a result, $\mu_{\lambda_{m}}=\{(z_{1},mn_{1}),(z_{2},mn_{2}),\dots,(z_{k},mn_{k})\}=m\cdot\mu_{\lambda}$ for all $m\geq 1$. Moreover, all the networks $G_{m}$ have isomorphic history trees. This is because, at every round, each agent in any of the copies of $V_{i}$ receives exactly two messages from other agents in copies of $V_{i}$ and exactly $m_{j}$ messages from agents in copies of $V_{j}$, for all $j\neq i$. Thus, it can be proved by induction that all agents in the copies of $V_{i}$ have isomorphic vistas, regardless of $m$.
Due to the fundamental theorem of history trees Theorem 3.1, all the agents with input $z_{i}$ must give the same output $\psi(z_{i},\mu_{\lambda})=\psi(z_{i},\mu_{\lambda_{m}})=\psi(z_{i},m\cdot\mu_{\lambda})$, regardless of $m$. Hence, by definition, only frequency-based functions can be computed in these networks. ∎
Observe that the previous proof does not rule out some special functions that are not strictly frequency-based, because they differ on $k$-tuples $(n_{1},n_{2},\dots,n_{k})$ where $n_{i}\leq 2$ for some $i$. However, the proof can be extended to include those functions as well, provided that we allow networks to be multigraphs. Indeed, for $n_{i}=1$, the cycle $C_{i}$ reduces to two self-loops on the unique agent in $V_{i}$; for $n_{i}=2$, the cycle $C_{i}$ reduces to two parallel edges between the two agents in $V_{i}$.
@@
Assume for a contradiction that there is such an algorithm $\mathcal{A}$. Let $\mathcal{G}$ be a static network consisting of three agents forming a cycle, and assign input 0 to all of them. If the agents execute $\mathcal{A}$, they eventually output the mean value 0 and terminate, say in $t$ rounds.
Now construct a static network $\mathcal{G}^{\prime}$ consisting of a cycle of $2t+2$ agents $p_{1}$, $p_{2}$, …, $p_{2t+2}$; assign input $1$ to $p_{1}$ and input 0 to all other agents. It is easy to see that, from round 0 to round $t$, the vista of the agent $p_{t+1}$ is isomorphic to the vista of any agent in $\mathcal{G}$. Therefore, if $p_{t+1}$ executes $\mathcal{A}$, it terminates in $t$ rounds with the incorrect output 0. Thus, $\mathcal{A}$ is incorrect. ∎